What is the pH of 250.0 mL of an aqueous solution containing 0.615 g of the strong acid trifluoromethane sulfonic acid (CF3SO3H)?


1
Expert's answer
2020-12-03T13:50:39-0500

CF3SO3H + H2O = [CF3SO3]- + H3O+  complete dissociation as a strong acid


n=m/M=0.615g / 150g/mol = 0.0041 mol

c(CF3SO3H)=[H+]=n/V=0.0041/0.25=0.0164 mol/L

pH=-lg[H+]=1.8


pH=1.8